RC meeting and Run Plan, Thursday Oct. 5, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary


  • After the RC meeting yesterday, Hunter continued to diagnose the loss of communication with the target software. A full summary of the work is available here: https://logbooks.jlab.org/entry/4192606
  • During this time he re-connected the cell RTDs which had been disconnected on Tuesday.
  • The nature of the work in fixing the target software required/caused the holding field to change several times such that the cell depolarized partially.


  • NMR: 19.8%
  • Not high enough for production per Gordon's 30% limit, so made a plan to run on H2 reference cell until polarization was acceptable.


  • Beam window cooling jets alarmed. Assisted the TO in attaching a new gas canister.
  • Reference cell purged and filled x3, filled to 147psi H2


  • Ref cell xsof ysof values set at no target position. BPM and IC look good, moving to ref cell position in preparation for beam on ref cell.
  • 10uA,20uA,30uA look good.
  • 5uA/5min Ramp to 45uA.


  • NMR: 33.9%
  • Proceeded to take 40uA on pol He3 target.


  • NMR: 36.7%
  • Pushed to 45uA on pol He3 target.
  • Production for the rest of OWL shift.

08:02 - Now

  • Ongoing MCC beam studies + PSS system fixes.
  • Hall swept and put into controlled access around 09:15am.
  • GEM Access: Holly + GEM experts
  • HODO Access: Gary
  • ECal Access: Simona + Don


  • BSY Trip, MCC say extra time on todays activities of order 2h.
  • While installing PSS chain bypass, machine was tripped and halls dropped to indetermined status -> restricted.

14:00- Now

  • Working with Techs, MCC, Radcon on re-establishing controlled access and eventually beam permit

Accelerator status

  • Beam studies ongoing atleast for the rest of day shift.
  • Earlier PSS system issues now fixed
  • There will be a maintenance day during Day shift on Wednesday 11/10

If you want to get into the hall, please communicate with the RC to get this organized!

Data collected :

  • Wednesday SWING: 0.133 C
  • Thursday OWL: 1.250 C
  • Thursday DAY: 0.000 C
  • TOTAL: 1.383 C

Hall A Subsystems Status


  • Magnets are off due to hall being in restricted/controlled access. Lasers and oven are still on, so the target is still actively polarising.


  • No detector issues today. HV have been off all day.

Hall access requests

  • Opportinistic Access Request: Colling Jet (Todd/Ellen - Whiteboard in CH) -- is this outdated?
    • If anyone wants to make an access, let the RC know!

Shift schedule

  • Monday 10/09 Swing Shift Needs a TO!
  • Please check when your shifts are!


Please sign up for any open shifts!

Short Term Run plan

  • Sweep hall to controlled
  • Wait for MCC to complete beam studies.
  • Perform full beam checkout
  • Standard production running at 45uA, with NMRs every 3 hours.

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)